A control premium of 30% yields a control value of $125 per share. The resulting discount for lack of control must be:

$96.15

b.

23.08%

c.

70.00%

d.

30.00%
Answer #1

Price without Control = 125/1.30 = $96.15

Lack of Control Discount = (96.15 - 125)/125

Lack of Control Discount = 23.08%